Price and efficiency variances. Sunshine Foods manufactures pumpkin scones. For January 2017,
it budgeted to purchase and use 14,750 pounds of pumpkin at $0.92 a pound. Actual purchases and usage
for January 2017 were 16,000 pounds at $0.85 a pound. Sunshine budgeted for 59,000 pumpkin scones.
Actual output was 59,200 pumpkin scones.
1. Compute the flexible-budget variance.
2. Compute the price and efficiency variances.
Answer #1

Question 1:

Flexible budget variance of a cost item is the difference between the value of that cost item in flexible budget and its actual value. Value of that cost item in flexible budget is nothing but the standard cost for actual output.

Standard material per unit of output= budgeted material/ budgeted output= 14750/59000= 0.25 pounds of pumpkin per scone

Standard rate= $0.92 per pound of pumpkin

Standard material cost per unit of output= Standard material per unit of output* Standard rate= 0.25*0.92= $0.23 per scone

Actual output= 59200 units

Standard cost for actual output= Standard material cost per unit of output* actual output= 0.23*59200= $ 13616

Actual cost= actual materials used* actual rate= 16000*0.85= $ 13600

Flexible budget variance= Standard cost for actual output- Actual cost= 13616 -13600= $16 favourable

Since the standard cost is higher than the actual cost, this situation is advantageous to the business and hence the variance is favourable.

Question 2:

Material price variance

Material price variance is the variance arising due to the difference in actual rate from the standard rate. The variance is favourable when the standard rate is higher than the actual rate and it is unfavourable when the opposite is the case.

Material price variance= (standard rate- actual rate) actual material used= (0.92-0.85)* 16000= $ 1120 favourable.

Material usage variance

Material usage variance measures the efficiency or inefficiency of the production department in utilising the materials purchased and used in the production. It is favourable when the actual materials used is lower than the standard material for actual output used and unfavourable when the opposite is the case.

Standard material per unit of output= 0.25 pounds of pumpkin per scone

Standard material used for actual output= Standard material per unit of output* actual output= 0.25*59200= 14800

Material usage variance= (standard materials for actual output- actual materials used) standard rate= (14800-16000) 0.92= $1104 unfavourable

The variance is unfavourable because the actual materials used is higher than the standard materials for the actual output.

Checking the answer:

Flexible budget variance= Price variance + usage variance

16 favourable= 1120 favourable + 1104 unfavourable= 16 favourable

Therefore, the variances computed are correct.